OpenStudy (anonymous):

Algebra 2 Help Please: What is the equation of the line in slope-intercept form that passes through the point (–2, 4) and is perpendicular to the line y = –2x + 4?

OpenStudy (anonymous):

Answer Choices: y = –1/2x – 5 y = –1/2x – 2 y = 1/2x + 2 y = 1/2x + 5

OpenStudy (anonymous):

For perpendicular lines, the product of their slope is -1

OpenStudy (anonymous):

So, m1 * m2 =-1

OpenStudy (anonymous):

does it help?